1. Give the relationship between each pair of molecules of the formula C5H6ClF. Your options are constitutional isomers, the same molecule, or geometric isomers. (10 points)0 Сінн H-C-c=ċ-ċ-C-H НЕ НН -0 Ķ Ķ Çi F H H-C-c-c=ċ-C-H I-O-I I-O-I I-O-I HH 0 a. A and B b. A and C C. A and D d. B and C e.
  2. For the following molecule (10 points) Number and name the longest chain ,Name and list all substituents (use IUPAC naming), Write the full, IUPAC name
Answer #1

CL4 H : I-U C-C=C-C-C-H .. - - H fH H Same molecule @ 6 A and BB A and c Same molecule. - H 4 Cl. FH C-c-c-c-C-lt 154 3 2 1 5cand D (constitutional Isomer) H HCL FH _CH HEC-C-C= I 1 + + H .. . lupac (2.Byomo, sb 4 Cyclo proyl, 5 blouro, 3, 1so propu!